### v1

# Batalin-Fradkin-Vilkovisky Quantization of FLPR model

Abstract: We quantize the Friedberg-Lee-Pang-Ren (FLPR) model within the framework of Batalin-Fradkin-Vilkovisky (BFV) formalism. We construct the nilpotent Becchi-Rouet-Stora-Tyutin (BRST) charges using constraints and the fermionic gauge-fixing function by means of admissible gauge conditions. We also derive the BRST invariant effective action and corresponding symmetries of the model in both polar and Cartesian coordinates. Additionally, with the help of BRST charge, we demonstrate that the physical states of the system are annihilated by the first-class constraints, which is consistent with the Dirac formalism.
PACS Nos: 11.15.-q, 11.10.Ef, 11.30.-j
Keywords: BFV quantization; FLPR model; Constrained systems; BRST symmetries.
1 Introduction
Constrained theories are the cornerstone of modern theoretical physics, as they include gauge theories, gravity theories, super-symmetric theories etc. The conventional methods of quantization are not directly applicable to such theories. Dirac developed a quantization technique to deal with such constrained systems - known as the Dirac formalism [1]. Later, a geometrically motivated approach based on the symplectic structure of the phase space - the Faddeev-Jackiw (FJ) formalism - was proposed [2]. The FJ formalism is also quite popular and widely used (see e.g. [3, 4] and references therein). As far as quantization of such constrained theories is concerned, the Becchi-Rouet-Stora-Tyutin (BRST) formalism is one of the most broadly used techniques [5, 6].
However, an another versatile method for the quantization of constrained system is introduced by Batalin, Fradkin and Vilkovisky and popularly known as the BFV formalism [7, 8]. It provides a systematic Hamiltonian framework for quantizing theories with first-class constraints which also preserves the BRST symmetries of the system. The BFV formalism has been generalized to include second-class theories by employing a conversion technique called the BFT formalism [9]. These methods have been explored in various classical and field theoretical models such as Chiral Schwinger model, Proca theory, QED and $(2+1)$-dimensional Hořava theory [10, 11, 12, 13].
The Friedberg-Lee-Pang-Ren (FLPR) model is a soluble gauge model that describes the motion of a single non-relativistic particle of unit mass governed by a central potential [14]. This is a $(0+1)$-dimensional quantum mechanical model with the characteristics of Gribov ambiguity, inspired by the dimensional reduction of Yang-Mills theory [15]. The BRST invariant path integral for the model is formulated by summing over all Gribov-type copies [16]. In addition, the FLPR model is being quantized using an admissible gauge-fixing condition within the framework of FJ quantization [17]. The system is also explored using a physical projector operator which avoids the gauge-fixing procedure [18]. Moreover, the existence of some novel off-shell nilpotent and absolutely anti-commuting symmetries of the FLPR model is established in [15]. Recently, this model is being investigated by means of supervariable and (anti-)chiral supervariable approaches [19, 21, 20].
The prime motive of our present endeavor is to quantize the FLPR model using an admissible gauge-fixing condition, in both polar and Cartesian coordinates, by employing BFV formalism. We intend to construct the nilpotent and conserved BRST charge and subsequently derive the BRST invariant gauge-fixed effective action and corresponding symmetries. In addition, we plan to show that the physical states of the theory are annihilated by the conserved BRST charge which is in accordance with the Dirac-quantization conditions.
The content of the present manuscript is organized as follows. In section 2, we provide the constraint structure of the FLPR model in polar coordinates using Dirac’s prescription of constraint analysis. The BFV quantization of the model is carried out in subsection 2.1. Our section 3 deals with the constraint analysis of the FLPR model in Cartesian coordinates whereas BFV formulation, in Cartesian coordinates, is presented in subsection 3.1. The section 4 discusses about conserved and nilpotent BRST charges and physicality criteria. Finally, we conclude our results in section 5.
2 FLPR model in polar coordinates: Constraint structure
We begin with the Lagrangian that describes the dynamics of the FLPR model in polar coordinates as [14]
${L}=\frac{1}{2}\dot{r}^{2}+\frac{1}{2}r^{2}(\dot{\theta}-g\zeta)^{2}+\frac{1}{2}(\dot{z}-\zeta)^{2}-V(r),$ (1)
where $\dot{r}$, $\dot{\theta}$ and $\dot{z}$ are generalized velocities corresponding to the generalized coordinates $r$, $\theta$ and $z$, respectively. Here, $\zeta$ represents a gauge variable and $g$ is coupling constant. The canonical momenta corresponding to the variables $r$, $\theta$, $z$ and $\zeta$ are given, respectively, as
$p_{r}=\dot{r},\qquad p_{\theta}=r^{2}(\dot{\theta}-g\zeta),\qquad p_{z}=\dot{z}-\zeta,\qquad p_{\zeta}=0.$ (2)
We identify $\Omega_{1}\equiv p_{\zeta}\approx 0$ as the primary constraint in the theory à la Dirac’s prescription for classification of constraints [1]. The canonical Hamiltonian $(H_{c})$ can be obtained as follows
${H}_{c}=\frac{1}{2}p_{r}^{2}+\frac{1}{2r^{2}}p_{\theta}^{2}+\frac{1}{2}p_{z}^{2}+\zeta(gp_{\theta}+p_{z})+V(r).$ (3)
The consistency condition on primary constraint $\Omega_{1}$ leads to a secondary constraint $(\Omega_{2})$
$\Omega_{2}\equiv gp_{\theta}+p_{z}\approx 0.$ (4)
There are no further constraints in the theory. Thus, the theory possess only two constraints, which are first-class in nature [1, 17].
2.1 BFV quantization
In order to make use of the BFV formalism, we introduce two sets of canonical ghost and anti-ghost variables $\big{(}{\cal{C}}^{i},{\cal\bar{P}}_{i}\big{)}$ and $\big{(}{\cal{P}}^{i},{\cal\bar{C}}_{i}\big{)}$ along with the auxiliary variables $\big{(}{N}^{i},{B}_{i}\big{)}$, such that they satisfy following super-Poisson bracket algebra
$\begin{gathered}\{{\cal{C}}^{i}(x),{\cal\bar{P}}_{j}(y)\}=\delta_{j}^{i}\delta(x-y),\\ \{{\cal{P}}^{i}(x),{\cal\bar{C}}_{j}(y)\}=\delta_{j}^{i}\delta(x-y),\\ \{{N}^{i}(x),{B}_{j}(y)\}=\delta_{j}^{i}\delta(x-y),\end{gathered}$ (5)
where $i,j=1,2$ [7, 8]. The super-Poisson bracket between any two variable $A$ and $B$ is defined as
$\{A,B\}=\frac{\partial_{r}A}{\partial q}\frac{\partial_{l}B}{\partial p}-(-1)^{\eta_{A}\eta_{B}}\frac{\partial_{r}B}{\partial q}\frac{\partial_{l}A}{\partial p},$ (6)
where $\eta_{A}$($\eta_{B}$) denote the number of fermions, called the ghost number, in A(B) and the subscript $r$ and $l$ denotes the right and left derivatives, respectively [10].
According to the BFV formalism, the nilpotent BRST charge $(Q)$ is constructed using constraints as follows
$Q={\cal{C}}^{1}\,p_{\zeta}+{\cal{C}}^{2}\,(gp_{\theta}+p_{z})+{\cal{P}}^{1}\,B_{1}+{\cal{P}}^{2}\,B_{2},$ (7)
and the gauge-fixing function $(\Phi)$ is defined as
$\Phi={\cal\bar{C}}_{1}\,\chi^{1}+{\cal\bar{C}}_{2}\,\chi^{2}+{\cal\bar{P}}_{1}\,N^{1}+{\cal\bar{P}}_{2}\,N^{2}.$ (8)
Here, we have chosen the admissible gauge-fixing conditions $\chi^{1}=\zeta$ and $\chi^{2}=z-\frac{1}{2}\,B_{2}.$ The minimal Hamiltonian $(H_{m})$ takes the following form
$H_{m}$ $=$ $H_{c}+\bar{{\cal{P}}_{a}}V^{a}_{b}{\cal{C}}^{b},$ (9)
$=$ $\frac{1}{2}p_{r}^{2}+\frac{1}{2r^{2}}p_{\theta}^{2}+\frac{1}{2}p_{z}^{2}+\zeta(gp_{\theta}+p_{z})+V(r)+{\cal\bar{P}}_{2}\,{\cal{C}}^{1},$
where we assume that the constraints fulfill the algebra: $\{H_{c},\Omega_{b}\}=\Omega_{a}V^{a}_{b}$. At this juncture, it is worth mentioning that the BRST charge $(Q)$ is conserved and fermionic in nature. As it is straightforward to verify that [11]
$\{Q,\;H_{m}\}=0,\qquad Q^{2}=\{Q,\;Q\}=0.$ (10)
In addition, the fermionic gauge-fixing function $(\Phi)$ and the BRST charge $(Q)$ satisfy the following algebra
$\{\{\Phi,\;Q\},\;Q\}=0.$ (11)
Now, the effective action in the extended phase space can be written as
$S_{eff}=\int dt\,\Big{[}p_{r}\,\dot{r}+p_{\theta}\,\dot{\theta}+p_{z}\,\dot{z}+p_{\zeta}\dot{\zeta}+B_{1}\,\dot{N^{1}}+B_{2}\,\dot{N^{2}}+{\cal\bar{P}}_{1}\,{\cal{\dot{C}}}^{1}+{\cal\bar{P}}_{2}\,{\cal{\dot{C}}}^{2}+{\cal\bar{C}}_{1}\,{\cal{\dot{P}}}^{1}+{\cal\bar{C}}_{2}\,{\cal{\dot{P}}}^{2}-H_{total}\Big{]},$ (12)
where $H_{total}=H_{m}+\{Q,\;\Phi\}$. We replace $\chi^{1}$ with $\chi^{1}+\dot{N}^{1}$ and subsequently omit the terms $\int dt\,(B_{1}\,\dot{N^{1}}\,+\,{\cal\bar{C}}_{1}\,{\cal{\dot{P}}}^{1}\,)\,$$\{Q,\int dt\,{\cal\bar{C}}_{1}\,\dot{N}^{1}\}$ in the above Eq. (12). Thus, the effective action takes the following form:
$S_{eff}=\int dt\,\Big{[}p_{r}\,\dot{r}+p_{\theta}\,\dot{\theta}+p_{z}\,\dot{z}+p_{\zeta}\dot{\zeta}+B_{2}\,\dot{N^{2}}+{\cal\bar{P}}_{1}\,{\cal{\dot{C}}}^{1}+{\cal\bar{P}}_{2}\,{\cal{\dot{C}}}^{2}++{\cal\bar{C}}_{2}\,{\cal{\dot{P}}}^{2}-H_{total}\Big{]}.$ (13)
In order to obtain the covariant effective action, we first eliminate the variables $B_{1}$, $N^{1}$, ${\cal\bar{C}}_{1}$, ${\cal{P}}^{1}$, ${\cal\bar{P}}_{1}$, ${\cal{C}}^{1}$, $\zeta$ and $p_{\zeta}$ by using Gaussian integration. Consequently, the explicit form of the effective action turns out to be
$S_{eff}$ $=$ $\int dt\,\Big{[}p_{r}\,\dot{r}+p_{\theta}\,\dot{\theta}+p_{z}\,\dot{z}+B\,\dot{N}+{\cal\bar{P}}\,{\cal{\dot{C}}}+{\cal\bar{C}}\,{\cal{\dot{P}}}-\frac{1}{2}\,p_{r}^{2}-\frac{1}{2r^{2}}\,p_{\theta}^{2}-\frac{1}{2}\,p_{z}^{2}-V(r)$ (14)
- $(gp_{\theta}+p_{z})\,N-{\cal\bar{C}}\,{\cal{C}}-B\Big{(}z-\frac{1}{2}\,B\Big{)}-{\cal\bar{P}}\,{\cal{P}}\Big{]},$
with $N^{2}\equiv N$, $B_{2}\equiv B$, ${\cal\bar{C}}_{2}\equiv{\cal\bar{C}}$, ${\cal{C}}^{2}\equiv{\cal{C}}$, ${\cal\bar{P}}_{2}\equiv{\cal\bar{P}}$ and ${\cal{P}}^{2}\equiv{\cal{P}}$. Second, the variation with respect to $p_{r}$, $p_{\theta}$, $p_{z}$, $\cal{P}$} and $\cal\bar{P}$} yields the following relations
$p_{r}=\dot{r},\qquad p_{\theta}=r^{2}\,(\dot{\theta}-gN),\qquad p_{z}=\dot{z}-N,\qquad{\cal\bar{P}}=-{\dot{\bar{\cal{C}}}},\qquad{\cal{P}}=\dot{\cal{C}}.$ (15)
Finally, substituting Eqs. (15) into the action (14) and identifying $N$ with $\zeta$ (cf. (2)), we obtain the covariant effective action
$S_{eff}=\int\text{d}t\,\Big{[}\frac{1}{2}\,\dot{r}^{2}+\frac{1}{2}\,r^{2}\big{(}\dot{\theta}-g\zeta\big{)}^{2}+\frac{1}{2}\big{(}\dot{z}-\zeta\big{)}^{2}-V(r)+B\Big{(}\dot{\zeta}-z+\frac{1}{2}\,B\Big{)}-{\dot{\bar{\cal{C}}}}\,{\dot{\cal{C}}}-{\cal{\bar{C}}}\,{\cal{C}}\Big{]}.$ (16)
This $(S_{eff})$ is invariant under the following off-shell nilpotent BRST symmetry transformations $(\delta_{b})$
$\begin{split}{\delta}_{b}r=0,\quad{\delta}_{b}\theta=g{\cal{C}},\quad{\delta}_{b}z={\cal{C}},\quad{\delta}_{b}\zeta=\dot{\cal{C}},\quad{\delta}_{b}{\cal{C}}=0,\quad{\delta}_{b}{\cal\bar{C}}=-B,\quad{\delta}_{b}B=0,\end{split}$ (17)
as $\delta_{b}L_{eff}=\frac{d}{dt}(B\dot{\cal{C}})$.
At this juncture, it should be noted that if we take variation over $\cal{C}$ and $\cal\bar{C}$ instead of $\cal{P}$ and $\cal\bar{P}$ in Eq. (14), we obtain the following effective action
$S^{\prime}_{eff}=\int\text{d}t\,\Big{[}\frac{1}{2}\,\dot{r}^{2}+\frac{1}{2}\,r^{2}\big{(}\dot{\theta}-g\zeta\big{)}^{2}+\frac{1}{2}\big{(}\dot{z}-\zeta\big{)}^{2}-V(r)+B\Big{(}\dot{\zeta}-z+\frac{1}{2}\,B\Big{)}-{\dot{\bar{\cal{P}}}}\,{\dot{\cal{P}}}-{\cal\bar{P}}\,{\cal{P}}\Big{]}.$ (18)
This form of effective action ($S^{\prime}_{eff}$) can be obtained by the variable transformation $\cal{C}\longrightarrow\cal{P}$ and $\cal\bar{C}\longrightarrow\cal\bar{P}$ in Eq. (16) and vice-versa.
3 FLPR model in Cartesian coordinates: Constraint structure
The Lagrangian that describes the FLPR model in Cartesian coordinates is given by [14, 18]
$\tilde{L}=\frac{1}{2}\Big{[}(\dot{x}+gy\zeta)^{2}+(\dot{y}-gx\zeta)^{2}+(\dot{z}-\zeta)^{2}\Big{]}-U(x^{2}+y^{2}),$ (19)
where $\dot{x}$, $\dot{y}$ and $\dot{z}$ represent the generalized velocities, $\zeta$ denotes the gauge variable and $g$ is a coupling constant. The canonical momenta corresponding to the coordinates $x$, $y$, $z$ and $\zeta$ are obtained as follows:
$p_{x}=\dot{x}+gy\zeta,\qquad p_{y}=\dot{y}-gx\zeta,\qquad p_{z}=\dot{z}-\zeta,\qquad p_{\zeta}=0.$ (20)
We identify $\tilde{\Omega}_{1}\equiv p_{\zeta}\approx 0$ as a primary constraint in the theory. The canonical Hamiltonian $({\tilde{H}_{c}})$ of the system is given as
$\tilde{H}_{c}=\frac{1}{2}\Big{(}p_{x}^{2}+p_{y}^{2}+p_{z}^{2}\Big{)}+\zeta\Big{(}g(xp_{y}-yp_{x})+p_{z}\Big{)}+U(x^{2}+y^{2}).$ (21)
The time evolution of the primary constraint yields a secondary constraint $(\tilde{\Omega}_{2})$ as follows:
$\tilde{\Omega}_{2}=g\,(xp_{y}-yp_{x})+p_{z}.$ (22)
There are only two constraints $(\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{2})$ in the theory and these are first-class in nature [1, 17].
3.1 BFV quantization
To employ BFV formalism, we introduce a pair of canonically conjugate ghost and anti-ghost variables along with their auxiliary variables as [7, 10]
$\big{(}{\cal{C}}^{i},{\cal\bar{P}}_{i}\big{)},\qquad\big{(}{\cal{P}}^{i},{\cal\bar{C}}_{i}\big{)},\qquad\big{(}{N}^{i},{B}_{i}\big{)},$ (23)
where $i=1,2.$ These variables satisfy the super-Poisson algebra given in Eq. (5). As per the BFV formulation, the nilpotent BRST charge ($\tilde{Q}$) and the fermionic gauge-fixing function ($\tilde{\Phi}$) are defined using constraints and gauge-fixing conditions, respectively, as follows:
$\tilde{Q}$ $=$ ${\cal{C}}^{1}\,p_{\zeta}+{\cal{C}}^{2}\,\big{(}g\,(xp_{y}-yp_{x})+p_{z}\big{)}+{\cal{P}}^{1}\,B_{1}+{\cal{P}}^{2}\,B_{2},$ (24)
$\tilde{\Phi}$ $=$ ${\cal\bar{C}}_{1}\,\tilde{\chi}^{1}+{\cal\bar{C}}_{2}\,\tilde{\chi}^{2}+{\cal\bar{P}}_{1}\,N^{1}+{\cal\bar{P}}_{2}\,N^{2},$ (25)
where $\tilde{\chi}^{1}=\zeta$ and $\tilde{\chi}^{2}=z-\frac{1}{2}B_{2}$ are the admissible gauge-fixing conditions. The minimal Hamiltonian $(\tilde{H}_{m})$ takes the following form
$\tilde{H}_{m}$ $=$ $\tilde{H}_{c}+\bar{{\cal{P}}_{a}}V^{a}_{b}{\cal{C}}^{b},$ (26)
$=$ $\frac{1}{2}\Big{(}p_{x}^{2}+p_{y}^{2}+p_{z}^{2}\Big{)}+\zeta\Big{(}g(xp_{y}-yp_{x})+p_{z}\Big{)}+U(x^{2}+y^{2})+{\cal\bar{P}}_{2}\,{\cal{C}}^{1},$
where we assume that the constraints fulfill the constraint algebra $\{H_{c},\Omega_{b}\}=\Omega_{a}V^{a}_{b}$. Moreover, $\tilde{Q}$, $\tilde{\Phi}$ and $\tilde{H}_{m}$ satisfy the algebra as listed in Eqs. (10) and (11).
Hence, the effective action in the extended phase space can be given as
$\tilde{S}_{eff}=\int\text{d}t\Big{[}p_{x}\dot{x}+p_{y}\dot{y}+p_{z}\dot{z}+p_{\zeta}\dot{\zeta}+B_{2}\dot{N^{2}}+{\cal\bar{P}}_{1}{\cal{\dot{C}}}^{1}+{\cal\bar{P}}_{2}{\cal{\dot{C}}}^{2}+{\cal\bar{C}}_{2}{\cal{\dot{P}}}^{2}\Big{]}-\tilde{H}_{total},$ (27)
where $\tilde{H}_{total}=\tilde{H}_{m}+\{\tilde{Q},\tilde{\Phi}\}$. In the above, the terms $\int dt\,(B_{1}\,\dot{N^{1}}+{\cal\bar{C}}_{1}\,{\cal{\dot{P}}}^{1}$ ) =  $\{\tilde{Q},\int dt\,\;{\cal\bar{C}}_{1}\,\dot{N}^{1}\}$ in Legendre transformations are suppressed by replacing $\tilde{\chi}^{1}$ with $\tilde{\chi}^{1}+\dot{N}^{1}$. Now, in order to obtain the effective action, first we perform the path integration over the variables $B_{1}$, $N^{1}$, ${\cal\bar{C}}_{1}$, ${\cal{P}}^{1}$, ${\cal\bar{P}}_{1}$, ${\cal{C}}^{1}$, $\zeta$ and $p_{\zeta}$ using Gaussian integration. Thus, we obtain
$\tilde{S}_{eff}$ $=$ $\int\text{d}t\Big{[}p_{x}\dot{x}+p_{y}\dot{y}+p_{z}\dot{z}+B\dot{N}+{\cal\bar{P}}{\cal{\dot{C}}}+{\cal\bar{C}}{\cal{\dot{P}}}-\frac{1}{2}p_{x}^{2}-\frac{1}{2}p_{y}^{2}-\frac{1}{2}p_{z}^{2}-U(x^{2}+y^{2})$ (28)
- $\big{(}g(xp_{y}-yp_{x})+p_{z}\big{)}N-B(z-\frac{1}{2}B)-{\cal\bar{C}}{\cal{C}}-{\cal\bar{P}}{\cal{P}}\Big{]},$
here we have redefined $N^{2}\equiv N$, $B_{2}\equiv B$, ${\cal\bar{C}}_{2}\equiv{\cal\bar{C}}$, ${\cal{C}}^{2}\equiv{\cal{C}}$, ${\cal\bar{P}}_{2}\equiv{\cal\bar{P}}$ and ${\cal{P}}^{2}\equiv{\cal{P}}$. Second, taking the variations over $p_{x}$, $p_{y}$, $p_{z}$, $\cal{P}$} and $\cal\bar{P}$} yields the following relations
$p_{x}=\dot{x}+gyN,\qquad p_{y}=\dot{y}-gxN,\qquad p_{z}=\dot{z}-N,\qquad{\cal\bar{P}}=-\dot{\bar{\cal{C}}},\qquad{\cal{P}}=\dot{\cal{C}}.$ (29)
Finally, identifying $N$ with $\zeta$ (cf. (20)) and substituting Eqs. (29) in action (28), we obtain the following BRST invariant effective action
$\tilde{S}_{eff}$ $=$ $\int\text{d}t\,\Big{[}\frac{1}{2}\big{(}(\dot{x}+gy\zeta)^{2}+(\dot{y}-gx\zeta)^{2}+(\dot{z}-\zeta)^{2}\big{)}+U(x^{2}+y^{2})$ (30)
$+$ $B(\dot{\zeta}-z+\frac{1}{2}B)-{\dot{\bar{\cal{C}}}}\dot{\cal{C}}-\bar{\cal{C}}{\cal{C}}\Big{]}.$
Corresponding BRST symmetry transformations $(\tilde{\delta}_{b})$ are listed below
$\begin{split}&\tilde{\delta}_{b}x=-gy{\cal{C}},\qquad\tilde{\delta}_{b}y=gx{\cal{C}},\qquad\tilde{\delta}_{b}z={\cal{C}},\qquad\tilde{\delta}_{b}\zeta=\dot{\cal{C}},\\ &\tilde{\delta}_{b}{\cal{C}}=0,\qquad\tilde{\delta}_{b}\bar{\cal{C}}=-B,\qquad\tilde{\delta}_{b}B=0,\end{split}$ (31)
which leaves the action $(\tilde{S}_{eff})$ invariant.
It is worthwhile to mention that, if we perform path integration over $\cal{C}$ and $\cal\bar{C}$ instead of $\cal{P}$} and $\bar{\cal{P}}$} in Eq. (28), the resulting effective action turns out to be
$\tilde{S}_{eff}$ $=$ $\int\text{d}t\,\Big{[}\frac{1}{2}\big{(}(\dot{x}+gy\zeta)^{2}+(\dot{y}-gx\zeta)^{2}+(\dot{z}-\zeta)^{2}\big{)}+U(x^{2}+y^{2})$ (32)
$+$ $B(\dot{\zeta}-z+\frac{1}{2}B)-{\dot{\bar{\cal{P}}}}\dot{\cal{P}}-\bar{\cal{P}}{\cal{P}}\Big{]}.$
This form of effective action can also obtained by the transformation $\cal{C}\longrightarrow\cal{P}$ and $\cal\bar{C}\longrightarrow\cal\bar{P}$ in Eq. (30) and vice-versa.
4 Physicality criteria: BRST charge
We note that, in both the polar and Cartesian coordinates, the conserved BRST charges
$Q={\cal{C}}\,(gp_{\theta}+p_{z})\,+\,\dot{\cal{C}}p_{\zeta},\qquad\tilde{Q}={\cal{C}}\big{(}g\,(xp_{y}-yp_{x})+p_{z}\big{)}+\dot{\cal{C}}p_{\zeta},$ (33)
are the generators of BRST symmetry transformations (cf. (17) and (31)) under which the effective actions $(S_{eff}$ and $\tilde{S}_{eff})$ remain invariant, respectively. In the above equation, we have used the Euler-Lagrange equations of motion ${\cal{P}}=\dot{\cal{C}}$ and the expression for momenta $p_{\zeta}=B$. These charges are also nilpotent of order two (cf. (10)).
In the BRST-quantized theory the total Hilbert space contains both the physical and ghost states as the direct product. Moreover, the ghost variables do not have any interactions with the physical variables of the system as they are decoupled. So, as the requirement of the physicality criteria, the dynamically stable subspace of states is annihilated by the above mentioned charges, i.e.;
[…]
which leads to the following conditions, respectively
[…]
$\big{(}g\,(xp_{y}-yp_{x})+p_{z}\big{)}\,|\tilde{\psi}\rangle=0,\qquad p_{\zeta}\,|\tilde{\psi}\rangle=0.$ (36)
It is important to note that $(gp_{\theta}+p_{z})\approx 0$ and $p_{\zeta}\approx 0$ are first-class constraints of the system in the polar coordinates whereas $\big{(}g\,(xp_{y}-yp_{x})+p_{z}\big{)}\,\approx 0$ and $p_{\zeta}\,\approx 0$ are first-class constraints in Cartesian coordinates. Therefore, physical states are annihilated by the first-class constraints, which is consistent with the results in [17] and Dirac formalism.
5 Conclusions
We have performed the BFV quantization of the FLPR model in both polar and Cartesian coordinates using admissible gauge conditions. The FLPR model is endowed with first-class constraints à la Dirac’s prescription for classification of constraints. In order to apply BFV formalism, we have introduced a pair of ghost and anti-ghost variables along with auxiliary variables. The conserved and nilpotent charge has been constructed using constraints. Whereas the fermionic gauge-fixing function is being put together by means of admissible gauge-fixing conditions. We have obtained the effective action, in the extended phase space, by the usual Legendre transformation. By eliminating a few variables using Gaussian integration, we have established the BRST invariant gauge-fixed effective action and corresponding BRST symmetries of the FLPR model in both polar and Cartesian coordinates. We have also indicated that the effective action has a discrete symmetry in the ghost sector. In addition, using the physicality criterion, we have shown that the physical states of the theory are annihilated by the first-class constraints, which is consistent with the Dirac formalism.
Acknowledgment: SG would like to dedicate this paper in the fond memories of Prof. Victor O. Rivelles who was a kind mentor.